OpenStudy (anonymous):

Solve the equation 5x = 25 Answer 3 5 7 9

OpenStudy (anonymous):

5x = 25 x = 5 5*5 = 25.

OpenStudy (ars22):

5

OpenStudy (ars22):

5x=5 x=25/5 x=5

jhonyy9 (jhonyy9):

x=25/5 x=5